# On Deepfake Voice Detection - It’s All in the Presentation

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Fig. 1: The sequence of steps transforming the voice signal is more elaborate in real-world application (telephone banking call center in this example) than simply using the source deepfake audio file created by TTS tool (a). Instead, the fraudster presents the deepfake voice through the phone using a loudspeaker or via direct injection technique. This audio is then transmitted through the telephony network, completing the presentation phase (b). In the task phase, we arrive at the real-world setting - a phone call takes place to engage the bank’s call center agent (c). The feedback arrow indicates that the fraudster uses the deepfake tool with the end task in mind, e.g. by creating conversational phrases. Each phase of this sequence introduces one or more distortions to the source deepfake audio signal
@@
Augmented neural vocoders / codecs augmentation 213,785 263,770 Public datasets (VoxCeleb 1&2, Fisher Spanish, Mixer6, Mixed Phone, RSR2015, SITW, SRE 16/18/19, Switchboard) processed with neural codecs and vocoders.
@@ 2 A Guide to Realistic Deepfake Data
To achieve the most effective performance of a deepfake detection system for the telephone banking scenario, the dataset creation process must go through the full process illustrated in Figure 1. Fraudsters would have sourced original recordings of their target and used these to generate deepfake audio clips, speaking phrases that will allow for financial crime ("I’d like to transfer some money"). The original deepfake recordings are then distorted by one or more factors through each phase in the sequence.
@@
Data Category 2 (Presented) - Presented audio datasets was built by processing raw data (as returned by TTS engines) through the presentation phase (Figure 1(b)), however using only a subset of the data described in the Base category. For the deepfake audio presentation, we used two different smartphones as calling devices, namely the Samsung Galaxy A12 and Redmi Note 8 Pro. For direct audio injection, we used two methods: digital injection through Bluetooth, and analog injection through a wired connection into the smartphone’s microphone input. As playback devices, we used an ESI Aktiv studio monitor and JBL Change Essential portable Bluetooth loudspeaker. All the recordings were done in a room with no special acoustic treatment. The entire process was automated, with calls going into a staged call center for this purpose. This yielded $\sim$70 thousand audio samples, which were only used for model training.
@@
Data Category 4 (Augmented) - vocoder-synthesized data is used only to augment training datasets, and was developed by processing bonafide human speech using neural vocoder [18] and codec [19] augmentation. i.e. "pseudo-spoof" data was created at scale by using the same components at the core of TTS algorithms. We used the HIFI-GAN [20], WaveGrad [21] and WaveNet [22] vocoders, as well as the Encodec [23] and Vocos [24] codecs.
@@ 3.3 Results
Because the majority of works in scientific literature focuses on one or very few benchmark(s), we don’t intend to compare accuracy on a per-dataset basis. Rather we offer the first performance analysis across a large swathe of known benchmarks and show how our methodology leads to systems that are competitive across most of them. The exceptions to this claim are the ASVspoof19 and ASVspoof21 LA results, where SOTA systems reported EER at 1% or less, likely due to shortcut learning; the nearly identical performance of our system for ASV21LA and ASV21LA-HT confirms that we do not take advantage of silences or other shortcuts, trading off a small increase in lab benchmark error for large gains in terms of real-world robustness. It is also worth highlighting that model accuracy is consistent when comparing the Pool error to the Average, another sign of good model generalization.

### v2

@@ 1 Introduction
Fig. 1: The sequence of steps transforming the voice signal is more elaborate in real-world application (telephone banking call center in this example) than simply using the source deepfake audio file created by the TTS tool (a). In (b), the presentation phase, the fraudster presents the deepfake voice through the phone using a loudspeaker or via direct injection technique, which is then transmitted through the telephony network. In the task phase, we arrive at the real-world setting - a phone call takes place to engage the bank’s call center agent (c). The feedback arrow indicates that the fraudster uses the deepfake tool with the end task in mind, e.g. by creating conversational phrases. Each phase of this sequence introduces one or more distortions to the source deepfake audio signal.
@@
Augmented neural vocoders / codecs augmentation 213,785 263,770 Public datasets (VoxCeleb 1&2, Fisher Spanish, Mixer6, Mixer Phone, RSR2015, SITW, SRE 16/18/19, Switchboard) processed with neural codecs and vocoders.
@@ 2 A Guide to Realistic Deepfake Data
To achieve the most effective performance of a deepfake detection system for the telephone banking scenario, the dataset creation process must go through the full process illustrated in Figure 1. Fraudsters would have sourced original recordings of their target and used these to generate deepfake audio clips, speaking phrases that will allow for financial crime (“I’d like to transfer some money”). The original deepfake recordings are then distorted by one or more factors through each phase in the sequence.
@@
Data Category 2 (Presented) - Presented audio datasets was built by processing raw data (as returned by TTS engines) through the presentation phase (Figure 1(b)), however using only a subset of the data described in the Base category. For the deepfake audio presentation, we used two different smartphones as calling devices, namely the Samsung Galaxy A12 and Redmi Note 8 Pro. For direct audio injection, we used two methods: digital injection through Bluetooth, and analog injection through a wired connection into the smartphone’s microphone input. As playback devices, we used an ESI Aktiv studio monitor and JBL Charge Essential portable Bluetooth loudspeaker. All the recordings were done in a room with no special acoustic treatment. The entire process was automated, with calls going into a staged call center for this purpose. This yielded $\sim$70 thousand audio samples, which were only used for model training.
@@
Data Category 4 (Augmented) - vocoder-synthesized data is only used to augment training datasets, done by processing bonafide human speech using neural vocoder [18] and codec [19] augmentation. i.e. “pseudo-spoof” waveform data was created at scale by using the same components at the core of TTS algorithms. We used the HIFI-GAN [20], WaveGrad [21] and WaveNet [22] vocoders, as well as the Encodec [23] and Vocos [24] codecs.
@@ 3.3 Results
Because the majority of works in scientific literature focuses on one or very few benchmark(s), we don’t intend to compare accuracy on a per-dataset basis. Rather we offer the first performance analysis across a large swath of known benchmarks and show how our methodology leads to systems that are competitive across most of them. The exceptions to this claim are the ASVspoof19 and ASVspoof21 LA results, where SOTA systems reported EER at 1% or less, likely due to shortcut learning; the nearly identical performance of our system for ASV21LA and ASV21LA-HT confirms that we do not take advantage of silences or other shortcuts, trading off a small increase in lab benchmark error for large gains in terms of real-world robustness. It is also worth highlighting that model accuracy is consistent when comparing the Pool error to the Average, another sign of good model generalization.